# A Multi-Stage Auto-Context Deep Learning Framework for Tissue and Nuclei Segmentation and Classification in H&E-Stained Histological Images of Advanced Melanoma

### v1

@@ Abstract
In this work, using the PUMA challenge dataset, we proposed a novel multi-stage deep learning approach by combining tissue and nuclei information in a unified framework based on the auto-context concept to perform segmentation and classification in histological images of melanoma. Through pre-training and further post-processing, our approach achieved second and first place rankings in the PUMA challenge, with average micro Dice tissue score and summed nuclei F1-score of 73.40% for Track 1 and 63.48% for Track 2, respectively. Our implementation for training and testing is available at:https://github.com/NimaTorbati/PumaSubmit
@@
Tissue Segmentation Nuclei Segmentation Medical Image Analysis Machine Learning Deep Learning Melanoma Computational Pathology.
11footnotetext: * Equal contributions22footnotetext: ** Equal contributions
@@ 1 Introduction
Skin cancer is one of the most common types of cancer, with melanoma being the most aggressive form [13]. Although many non-invasive approaches such as dermatoscopic or clinical image analysis are widely used for early detection [21, 20, 18], examination through biopsy and analysis of histological images remains the gold standard for diagnosis [4].
@@
Various staining methods exist for histological specimens, but hematoxylin and eosin (H&E) staining is the most commonly used by pathologists [8]. To categorize histological images, such as distinguishing between primary melanomas (referred to as primary in the rest of the text) and melanomas at distant metastatic sites (referred to as metastatic in the rest of the text), medical experts and pathologists can visually examine the slides and, based on their knowledge and experience, make diagnostic and treatment decisions. Several factors influence diagnostic accuracy, including the location and types of tissues and cells in the investigated specimens. Manual examination and assessment are labor-intensive and subject to intra- and inter-observer variability [28].
@@
With the advancement of machine learning and deep learning, automated computer-based methods have become powerful tools for performing tissue and nuclei segmentation and classification [19]. These models can be trained using various approaches, including unsupervised, semi-supervised, self-supervised, and recently developed contrastive learning methods used in foundation model training [22, 23, 16]. However, the most accurate models are typically trained in a supervised manner, which requires annotated data.
@@
For supervised training of tissue and nuclei segmentation and classification models, several publicly available datasets exist. Datasets such as BACH [2], GlaS [29], and CRAG [1] can be used for tissue segmentation. For nuclei segmentation and classification, datasets like PanNuke [5], NuInsSeg [15], MoNuSAC [32], and MoNuSeg [12] can be utilized.
@@
In terms of model development, many approaches have been proposed for each task. Convolutional neural networks (CNNs)-based and vision transformer (ViT)-based models such as U-Net [24], U-Net++ [35], SegFormer [33, 27], and nnU-Net [11, 30] have been used for tissue segmentation in histological images. On the other hand, HoverNet [7], HoVer-Next [3], CellViT [10], CellViT++ [9], and DDU-Net [17] are examples of models developed specifically for nuclei instance segmentation and classification.
While these datasets and models have contributed significantly to the field of computational pathology, only a few focus on the combined task of simultaneous tissue and nuclei segmentation and classification. Examples of such efforts include the OCELOT challenge dataset [26] and the models proposed in [6, 26]. However, these datasets and models are not specifically designed for histological image analysis of advanced melanoma.
To address this gap, the Panoptic segmentation of nUclei and tissue in advanced MelanomA (PUMA) challenge introduced a dataset specifically designed for simultaneous tissue and nuclei segmentation and classification in H&E-stained melanoma specimens. The challenge aims to provide a unique dataset and encourage the research community to develop robust models for two main tasks defined in its two tracks. Track 1 involves segmentation and classification of five tissue types and three nuclei instance types. Track 2 includes segmentation and classification of the same five tissue types and ten nuclei instance classes. It is important to note that the background class is not treated as a separate class in either track.
@@
In this study, we proposed a novel multi-stage approach based on the auto-context concept [31] and participated in both tracks of the PUMA challenge. The results on the final test set confirmed the excellent performance of our method, achieving a mean micro Dice tissue score and summed nuclei F1-score of 73.40% for Track 1 and 63.48% for Track 2, respectively. These results ranked our approach (submitted under the name "LSM team") second in Track 1 and first in Track 2 of the PUMA competition, respectively  [^1] https://puma.grand-challenge.org/.
@@ 2 Method
We proposed a multi-stage auto-context pipeline, where the outputs of earlier stages are used as contextual input to refine tissue segmentation and nuclei segmentation and classification in later stages. Our method consists of four main stages, as illustrated in Figure 1, Figure 2, Figure 3, and Figure 4.
In the first stage (Figure 1), we designed a classifier to determine the input image type, distinguishing between primary and metastatic frames. In the second stage (Figure 2), initial tissue segmentation is performed. The third stage (Figure 3) involves nuclei instance segmentation using the baseline HoVer-Next model, along with nuclei classification, which utilizes the tissue information from the previous step. Finally, in the fourth stage (Figure 4), tissue segmentation is refined using the nuclei masks produced in the third stage. Further explanation of our entire pipeline is described below.
@@ 2.1 Datasets
We used two datasets in this work: the PanopTILs dataset [14] and the PUMA dataset [28]. The former was used only for pre-training, while the latter was used for fine-tuning and participation in the PUMA challenge.
The PanopTILs dataset contains a total of 1,709 image patches extracted from whole-slide images. Each patch has a size of 1024×1024 pixels with an image resolution of 0.25 microns per pixel, which is approximately equivalent to 40× magnification. Since the PUMA dataset images have a slightly different resolution (0.23 microns per pixel), we first cropped the PanopTILs images to 880$\times$880 pixels (top-left cropping) and then resized them to 1024$\times$1024 pixels to ensure a similar resolution comparable to that of the PUMA images. The PanopTILs dataset includes annotations for 9 tissue types (exclude, cancerous epithelium, stroma, TILs, normal epithelium, junk/debris (necrosis), blood, other, and white space/empty) and 10 nuclei types (exclude, cancer nucleus, stromal nucleus, large stromal nucleus, lymphocyte nucleus, plasma cell / large TIL nucleus, normal epithelial nucleus, other nucleus, unknown/ambiguous nucleus, and background). From the original 1,709 image regions, we used only 300 images for pre-training and merged some classes to align the labels more closely with those used in the PUMA dataset. In the case of tissue classes, we grouped "exclude," "other," and "white space/empty" as white background, while "cancerous epithelium" and "TILs" were grouped as tumor. Only images containing necrosis or tumor tissue were used in our analysis, while those containing the normal epithelium class were excluded.
@@
The other dataset used in this study is the PUMA dataset [28], which served as the main dataset for our experiments. It contains a total of 310 melanoma image patches: 155 from primary tumors and 155 from metastatic tumors, all with a size of 1024$\times$1024 pixels. These were divided into training (206 images), validation (10 images), and test (94 images) sets. While both tissue and nuclei annotations were provided for the training images, the labels for the validation and test images were kept private by the challenge organizers. Participants in the challenge could evaluate their models on the 10 validation images (also referred as preliminary test set in the challenge) through the PUMA Grand Challenge platform  [^2] https://puma.grand-challenge.org/ and submit their final results for the 94 test images. The PUMA dataset includes annotations for six tissue types: tumor, stroma, epidermis, necrosis, blood vessel, and background. For nuclei, Track 1 of the challenge includes four types: tumor cells, lymphocyte, stroma cell, and other. Track 2 includes ten types: tumor cell, lymphocyte, plasma cell, histocyte, melanophage, neutrophil, stroma cell, endothelium, epithelium, and apoptotic cell, as well as the background class. In our experiments, we used all training images except one due to a mismatch between the image and the provided segmentation mask. Further details about the PUMA dataset can be found in [28].
@@
It is worth mentioning that the flowcharts shown in the rest of the manuscript present only the fine-tuning steps with the PUMA dataset and do not include the pre-training with the PanopTILs dataset.
@@ 2.3 Stage 1: Classifier
To address the challenges posed by the limited dataset size and variations in tissue types across skin and other organs, we differentiated tissue types between the primary and metastatic phases. Our approach involves training two separate models: one for primary images and another for metastatic images. Therefore, an initial classification step is required to determine the image type. To build the classifier, we first increased the number of tissue classes from five to ten, with the first five representing primary tissue types and the remaining five corresponding to metastatic tissue types. We then trained a SegFormer-B2 segmentation model with 11 classes (including the background) as the main classifier, followed by a set of handcrafted classification rules to categorize images as either primary or metastatic, as stated below.
An image is classified as primary if the segmentation result includes Epidermis. If the Epidermis is not present, the image is still classified as primary if the total number of primary pixels exceeds the total number of metastatic pixels. Otherwise, the image is classified as metastatic.
@@
Figure 1: The first stage of the proposed method: In the first stage of the proposed method, a SegFormer model is trained to classify the input image type (primary or metastatic) based on segmentation result and the classification rules.
@@ 2.4 Stage2: Initial Tissue Segmentation
During our experiments, we observed that the performance of the SegFormer model in detecting blood vessel tissue was suboptimal. To address this, we trained a U-Net model using all available images in the PUMA dataset specifically for blood vessel detection. As a result, in our approach, we replaced the SegFormer predictions for blood vessel tissue with the corresponding U-Net predictions.
@@
For blood vessel class, we directly use the U-Net output:
@@
Figure 2: The second stage of the proposed method: Initial tissue segmentation is performed using two SegFormer models, one for primary tissues and the other for metastatic tissues. Due to the Segformer model’s poor performance in segmenting blood vessel tissues, a U-Net model is trained separately for blood vessel detection and the results are refined using the tissue ensemble rules.
@@ 2.5 Stage3: Nuclei Segmentation
As observed in our experiments and inferred from the results in [28], rather than improving the instance segmentation outputs, we opted to use the baseline instance segmentation results from the HoVer-NeXt model and focused on enhancing the classification results using external models. Specifically, we employed U-Net++ with four input channels for this task. We previously tested this approach (that is, using an external model solely for nuclei classification) in the MoNuSAC challenge, where we achieved first place on the post-challenge leaderboard [17, 32]. In this study, for each detected instance from the HoVer-NeXt model, we determined the instance class from the U-Net++ predictions using a simple majority voting approach. The number of U-Net++ output classes depends on the challenge track: four classes for Track 1 and eleven classes for Track 2, including the background in both cases.
Figure 3: The third stage of the proposed method consists of two branches for nuclei segmentation, utilizing tissue segmentation information. In the first branch, a U-Net++ model is trained for nuclei class map detection, incorporating tissue segmentation results as an additional input channel. In the second branch, nuclei instance segmentation is performed using the HoVer-NeXt model. The final nuclei segmentation and classification is obtained by combining the class maps and instance masks through a majority voting approach.

### v2

@@ Abstract
In this work, using the PUMA challenge dataset, we propose a novel multi-stage deep learning approach by combining tissue and nuclei information in a unified framework based on the auto-context concept to perform segmentation and classification in histological images of melanoma. Through pre-training and further post-processing, our approach achieved second and first place rankings in the PUMA challenge, with average micro Dice tissue score and summed nuclei F1-score of 73.40% for Track 1 and 63.48% for Track 2, respectively. Furthermore, through a comprehensive ablation study and additional evaluation on an external dataset, we demonstrated the effectiveness of the framework components as well as the generalization capabilities of the proposed approach. Our implementation for training and testing is available at:https://github.com/NimaTorbati/PumaSubmit
@@
Tissue Segmentation , Nuclei Segmentation , Medical Image Analysis , Machine Learning , Deep Learning , Melanoma , Computational Pathology.
[…]
[label1]organization=Research Center for Medical Image Analysis and Artificial Intelligence, Department of Medicine, Faculty of Medicine and Dentistry, Danube Private University, city=Krems an der Donau, postcode=3500, country=Austria
[…]
[label2]organization=Department of Pathophysiology and Allergy Research, Center of Pathophysiology, Infectiology and Immunology, Medical University of Vienna, city=Vienna, postcode=1090, country=Austria
[…]
[label3]organization=Comprehensive Center for AI in Medicine (CAIM), Medical University of Vienna, city=Vienna, postcode=1090, country=Austria
[…]
[label4]organization=Research Center for Clinical AI-Research in Omics and Medical Data Science, Department of Medicine, Faculty of Medicine and Dentistry, Danube Private University, city=Krems an der Donau, postcode=3500, country=Austria
[…]
[label5]organization=Austrian Center for Medical Innovation and Technology, city=Wiener Neustadt, postcode=2700, country=Austria
@@ 1 Introduction
Skin cancer is one of the most common types of cancer, with melanoma being the most aggressive form [1]. Although many non-invasive approaches such as dermatoscopic or clinical image analysis are widely used for early detection [2, 3, 4], examination through biopsy and subsequent analysis of histological images remains the gold standard for diagnosis [5].
@@
Various staining methods exist for histological specimens, but hematoxylin and eosin (H&E) staining is the most commonly used by pathologists [6]. To distinguish between primary melanomas (referred to as primary in the rest of the text) and distant metastases arising from melanomas (referred to as metastatic in the rest of the text), for example, medical experts and pathologists visually examine H&E-stained slides and, based on their knowledge and experience, make diagnostic calls that guide treatment decisions. Several factors influence diagnostic accuracy, including the complexity of tissue anatomy and the cell type composition in the investigated specimens. Microscopic examination and assessment of histological tissue specimens by pathologist are labor-intensive and subject to intra- and inter-observer variability [7].
@@
With the advancement of machine learning and deep learning (DL), automated computer-based methods have become powerful tools for performing tissue and nuclei segmentation and classification [8]. These models can be trained using various approaches, including unsupervised, semi-supervised, self-supervised, and recently developed contrastive learning methods used in foundation model training [9, 10, 11]. However, the most accurate models are typically trained in a supervised manner, which requires annotated data.
@@
For supervised training of tissue and nuclei segmentation and classification models, several publicly available datasets exist. Datasets such as BACH [12], GlaS [13], and CRAG [14] can be used for tissue segmentation. For nuclei segmentation and classification, datasets like PanNuke [15], NuInsSeg [16], CryoNuSeg [17], MoNuSAC [18], and MoNuSeg [19] can be utilized.
@@
In terms of model development, various approaches have been proposed for each task. Convolutional neural networks (CNNs)-based and vision transformer (ViT)-based models such as U-Net [20], U-Net++ [21], SegFormer [22, 23], and nnU-Net [24, 25] have been used for tissue segmentation in histological images. On the other hand, HoverNet [26], HoVer-NeXt [27], CellViT [28], CellViT++ [29], and DDU-Net [30, 31] are examples of models developed specifically for nuclei instance segmentation and classification.
While these datasets and models have contributed significantly to the field of computational pathology, most of them treat tissue and nuclei segmentation as independent tasks, which may be suboptimal. Understanding the relationship between tissue context and nuclear characteristics has been shown to improve model performance for nuclei and tissue segmentation tasks [32, 33]. Furthermore, in specific clinical applications such as tumor-infiltrating lymphocyte (TIL) analysis, medical experts examine both tissue and nuclear features to support various diagnostic and prognostic assessments [34, 33, 35, 36, 37].
The developed models for joint segmentation of tissue and nuclei typically consist of two components: one for tissue and one for nuclei. These components can be configured to exchange information either in parallel [34, 33] or in series [38, 39, 40]. In methods with a parallel structure, the tissue and nuclei networks are trained simultaneously, and information fusion between them can be achieved through parallel encoders with interconnections [41], parallel decoders [42], or parallel models utilizing specialized skip connections within the decoder [33]. In methods with a serial structure, a model such as SegFormer [22], DeepLabv3+ [43], or a ViT-based model [39] is first trained for the tissue segmentation task. The obtained results, together with the original input image, are then used to perform nuclei segmentation. To train such models, a few publicly available datasets for various tissue types, such as TIGER, PanopTILs, OCELOT, and the recently published PUMA dataset, can be used. While these models have shown improved performance for tissue and nuclei segmentation tasks, further improvements are still needed. Moreover, their application to histological image analysis of advanced melanoma has not yet been investigated.
For DL-based melanoma histological image analysis, most developed methods focus on distinct tasks of nuclei and tissue segmentation [44], and only a few approaches have attempted to integrate both nuclei- and tissue-level information. Liu et al. [45] investigated the performance of the Mask R-CNN network [46, 47] for detecting potential melanoma regions using a sparse and noisy dataset annotated with both tissue and nuclei labels. However, training a model on such data without a systematic strategy for integrating tissue and nuclei information may be suboptimal. In another study [48], a method was proposed for melanoma region detection based solely on nuclei segmentation. This approach first employs a CNN-based model to perform nuclei segmentation, followed by binary morphological operations to group nuclei into melanoma regions. However, relying solely on nuclei-level features limits the model’s ability to capture global contextual dependencies. Moreover, the performance of such binary approaches is often sensitive to hyperparameters, such as the size of the dilation kernel. In [49], a weakly supervised framework was introduced to differentiate skin melanocytes from keratinocytes. This approach initially trains a DeepLabv3+ network to segment epidermal tissues. Patches extracted from these regions are then processed using a pre-trained VGG16-based classifier [50]. To differentiate melanocytes from keratinocytes, the method applies Otsu thresholding [51] to the gradient maps of the final classifier layer (similar to Grad-CAM [52]). While this method avoids the need for nuclei annotations, it is not readily applicable to multi-class or instance segmentation tasks, limiting its scalability and generalizability.
Considering the current limitations of existing models for joint nuclei and tissue segmentation in histological images of advanced melanoma, there is a clear need for a robust and generalizable framework to address these tasks. Another major limitation is that most of these models have been developed using private datasets, which restricts their reproducibility and broader applicability. To address the latter issue, the Panoptic segmentation of nUclei and tissue in advanced MelanomA (PUMA) challenge dataset [7] was recently introduced. This dataset contains expertly annotated images of both tissue and nuclei, and it may serve as a valuable benchmark for research specifically focused on melanoma histological image analysis. The PUMA challenge introduces a unique dataset and encourages the research community to develop robust models for two main tasks defined in its two tracks. Track 1 involves segmentation and classification of five tissue types and three nuclei instance types. Track 2 includes segmentation and classification of the same five tissue types and ten nuclei instance classes.
@@
In this study, using the PUMA challenge dataset, we propose a novel multi-stage approach based on the auto-context concept [53]. The proposed method consists of four sequential stages: classification, initial tissue segmentation, nuclei segmentation, and final tissue segmentation. In summary, in stage 1, a classifier is designed to enhance model performance by leveraging the biological characteristics of tissue structures. In stage 2, global contextual information is extracted through the segmentation of tissues. In stage 3, the tissue segmentation output from stage 2 is fused with the input image to guide nuclei segmentation. Finally, in stage 4, the nuclei segmentation result from stage 3 is fused with the input image to perform the final tissue segmentation. The results on the final test set of both tracks of the challenge confirmed the excellent performance of our method, achieving a mean micro Dice tissue score and summed nuclei F1-score of 73.40% for Track 1 and 63.48% for Track 2, respectively. These results ranked our approach (submitted under the name ”LSM team”) second in Track 1 and first in Track 2 of the PUMA competition, respectively  [^5] https://puma.grand-challenge.org/. Through a comprehensive ablation study, we demonstrated the effectiveness of each stage of our proposed model on overall performance. Additionally, we evaluated the performance of our approach for nuclei instance segmentation and classification on an external dataset (the MoNuSAC challenge data), and showed improved results compared to the top performers of the challenge.
@@ 2 Method
We propose a multi-stage auto-context pipeline, where the outputs of earlier stages are used as contextual input to refine tissue segmentation and nuclei segmentation and classification in later stages. Our method consists of four main stages, as illustrated in Figure 1, Figure 2, Figure 3, and Figure 4.
In the first stage (Figure 1), we designed a classifier to determine the input image type, distinguishing between primary and metastatic frames. In the second stage (Figure 2), initial tissue segmentation is performed. The third stage (Figure 3) involves nuclei instance segmentation using the baseline HoVer-NeXt model, along with nuclei classification, which utilizes the tissue information from the previous step. Finally, in the fourth stage (Figure 4), tissue segmentation is refined using the nuclei masks produced in the third stage. Further description of our pipeline is given below.
@@ 2.1 Datasets
We used three datasets in this work: the PanopTILs dataset [33], the PUMA dataset [7], and the MoNuSAC dataset [54], each for different purposes.
The PanopTILs dataset was used exclusively for pre-training. The PUMA dataset served as the main dataset for training and evaluation, while the MoNuSAC dataset was used for external validation of our proposed approach.
[…]
@@
The second dataset used in this study is the PUMA dataset, which served as the main dataset for our experiments. It contains a total of 310 melanoma image patches: 155 from primary tumors and 155 from metastatic tumors, all with a size of $\times1024$ pixels. These were divided into training (206 images), validation (10 images), and test (94 images) sets. While both tissue and nuclei annotations were provided for the training images, the corresponding labels for the validation and test sets were not disclosed by the challenge organizers. Participants in the challenge could evaluate their models on the 10 validation images (also referred as preliminary test set in the challenge) through the PUMA Grand Challenge platform  [^6] https://puma.grand-challenge.org/ and submit their final results for the 94 test images. The PUMA dataset includes annotations for six tissue types: tumor, stroma, epidermis, necrosis, blood vessel, and background. For nuclei, Track 1 of the challenge includes three types: tumor cells, lymphocyte, and other. Track 2 includes ten types: tumor cell, lymphocyte, plasma cell, histocyte, melanophage, neutrophil, stroma cell, endothelium, epithelium, and apoptotic cell, as well as the background class. In our experiments, we used all training images except one due to a mismatch between the image and the provided segmentation mask. Further details about the PUMA dataset can be found in [7].
As it is essential to demonstrate the generalization capability of the proposed pipeline [55], we used the MoNuSAC  [^7] https://monusac-2020.grand-challenge.org/ dataset as an external validation set. The MoNuSAC dataset comprises 310 image patches of varying sizes from H&E-stained tissue sections. Specimens were derived from patients with lung, prostate, kidney, and breast cancer. The dataset consists of 209 training images and 101 test images. It includes annotations for four types of nuclei: epithelial cells, lymphocytes, macrophages, and neutrophils. However, tissue-level annotations are not provided.
@@
It is worth noting that the flowcharts presented in the remainder of the manuscript illustrate only the fine-tuning steps using the PUMA dataset and do not depict the pre-training conducted with the PanopTILs dataset. The training and evaluation procedure for the MoNuSAC dataset is described separately in Section 2.8.
@@ 2.3 Stage 1: Classifier
To address the challenges posed by the limited dataset size and variations in tissue types, we differentiated tissue types between the primary tumors and metastatic sites. Our approach involves training of two separate models: one for primary images and another for metastatic images. Therefore, an initial classification step is required to determine the image type. To build the classifier, we first increased the number of tissue classes from five to ten, with the first five representing primary tissue types and the remaining five corresponding to metastatic tissue types. We then trained a SegFormer-B2 segmentation model with 11 classes (including the background) as the main classifier, followed by a set of handcrafted classification rules to categorize images as either primary or metastatic, as stated below.
Since epidermis is a skin-specific tissue, an image is classified as primary if the segmentation result includes ”epidermis” class. If the epidermis is not present, the image is still classified as primary if the total number of primary pixels exceeds the total number of metastatic pixels. Otherwise, the image is classified as metastatic.
@@
Figure 1: The first stage of the proposed method. In this stage, a SegFormer model is trained to classify the input image type (primary or metastatic) based on segmentation result and the classification rules.
@@ 2.4 Stage 2: Initial Tissue Segmentation
During our experiments, we observed that the performance of the SegFormer model in detecting blood vessels was suboptimal. To address this, we trained a U-Net model using all available training images in the PUMA dataset specifically for blood vessel detection. As a result, in our approach, we replaced the SegFormer predictions for blood vessels with the corresponding U-Net predictions.
@@
For the blood vessel class, as described, we directly use the U-Net output.
@@
The flowchart of this stage is shown in Figure 2.
Figure 2: The second stage of the proposed method. Initial tissue segmentation is performed using two SegFormer models, one for primary tissues and the other for metastatic tissues. Due to the Segformer model’s poor performance in segmenting blood vessels, a U-Net model is trained separately for blood vessel detection and the results are refined using the tissue ensemble rules. Color code for tissue segmentation: red for tumor, blue for stroma, and green for blood vessel.
@@ 2.5 Stage 3: Nuclei Segmentation
As observed in our experiments and inferred from the results in [7], rather than improving the instance segmentation outputs, we chose to use the baseline instance segmentation results from the HoVer-NeXt model and focused on enhancing the classification results using an external model. Specifically, we employed U-Net++ with four input channels for this task. We previously tested this approach (that is, using an external model solely for nuclei classification) in the MoNuSAC challenge, where we achieved first place on the post-challenge leaderboard [30, 18]. In this study, for each detected instance from the HoVer-NeXt model, we determined the instance class from the U-Net++ predictions using a simple majority voting approach. The number of U-Net++ output classes depends on the challenge track: four classes for Track 1 and eleven classes for Track 2, including the background in both cases.
Figure 3: The third stage of the proposed method. This stage consists of two branches, with the upper part dedicated to nuclei classification and the lower part to nuclei instance segmentation. In the first branch, a U-Net++ model is trained for nuclei class map detection, incorporating tissue segmentation results as an additional input channel. In the second branch, nuclei instance segmentation is performed using the HoVer-NeXt model. The final nuclei segmentation and classification is obtained by combining the class maps and instance masks through a majority voting approach. Color code for tissue segmentation: red for tumor, blue for stroma, and green for blood vessel; color code for nuclei segmentation: various cell types are indicated by different colors.